A special vehicle for traveling on glacial ice has tires that are 4.8 ft in diameter. If the vehicle travels at 3.5 mi/h, what is the angular velocity (in r/min) of the tire?
Answer by lwsshak3(6472)

A special vehicle for traveling on glacial ice has tires that are 4.8 ft in diameter. If the vehicle travels at 3.5 mi/h, what is the angular velocity (in r/min) of the tire?
..
Using quantitative analysis method:

(3.5 mi/hr)*(rev/4.8π ft)*(2π rad/rev*(5280 ft/mi)*(hr/60min)
all units and π cancel out except rad/min
=(3.5*2*5280)/(4.8*60) rad/min
=128.33 rad/min (ans)
..
factors (ratios) used:
rev/circumference=rev/π*diameter=rev/4.8π ft
2π rad/rev
5280 ft/mi
hr/60 min